OpenStudy (anonymous):

Find the number guaranteed by the Mean Value Theorem for the function f(x)=e^(0.5*x) on the interval [0,2] A) 1.083 B) 0.709 C) 0.614 D) -0.304 I keep getting 0.859 and don't know what I'm doing wrong

OpenStudy (anonymous):

This is what I'm doing: \[\frac{ e ^{\frac{ 1 }{ 2}*2} -e ^{\frac{ 1 }{ 2 }*2}{ }}{2-0 }\]

OpenStudy (anonymous):

And I get 0.859 for that :/